1. how many grams of the solid can be dissolved in 25g of water at 40c?
  2. assuming that at 40 c, the volume of the solution is equal to the volume of the water, what is the density of the solution formed above? (dh2o=1.00g/ml)
  3. a student adds 35g of the substance to 72g of water at 25 c. will it all dissolve? if not, how much remains undissolved?
1
Expert's answer
2021-01-28T05:36:45-0500

Using the solubility graph, grams of a solid dissolved in 25g of water at 25c is 62g

Therefore,

"=\dfrac{62*25}{ 100}=26.25"

26.25grams of a solid.

A saturated solution is a mixture of liquids that contains a maximum number of solutes that is capable of dissolving.

"35g+72g=107g"

Some will dissolve but some will not."7g\ will\ remain\ undissolved"